# DIFFender: Diffusion-Based Adversarial Defense against Patch Attacks

### v3

@@ Abstract
Adversarial attacks, particularly patch attacks, pose significant threats to the robustness and reliability of deep learning models. Developing reliable defenses against patch attacks is crucial for real-world applications, yet current research in this area is unsatisfactory. In this paper, we propose DIFFender, a novel defense method that leverages a text-guided diffusion model to defend against adversarial patches. DIFFender includes two main stages: patch localization and patch restoration. In the localization stage, we find and exploit an intriguing property of the diffusion model to precisely identify the locations of adversarial patches. In the restoration stage, we employ the diffusion model to reconstruct the adversarial regions in the images while preserving the integrity of the visual content. Thanks to the former finding, these two stages can be simultaneously guided by a unified diffusion model. Thus, we can utilize the close interaction between them to improve the whole defense performance. Moreover, we propose a few-shot prompt-tuning algorithm to fine-tune the diffusion model, enabling the pre-trained diffusion model to adapt to the defense task easily. We conduct extensive experiments on image classification, face recognition, and further in the physical world, demonstrating that our proposed method exhibits superior robustness under strong adaptive attacks and generalizes well across various scenarios, diverse classifiers, and multiple patch attack methods.
@@ 1 Introduction
Deep neural networks are vulnerable to adversarial examples [29, 9], in which imperceptible perturbations are intentionally added to natural examples, leading to incorrect predictions with high confidence of the model. Most adversarial attacks and defenses are devoted to studying the $\ell_{p}$-norm threat models [9, 3, 6, 21], which assume that the adversarial perturbations are restricted by the $\ell_{p}$ norm to be imperceptible. However, the classic $\ell_{p}$ perturbations require modification of every pixel of the images, which is typically not practical in the physical world. On the other hand, adversarial patch attacks [2, 17, 18, 33], which usually apply perturbations to a localized region of the objects, are more physically realizable. Adversarial patch attacks pose significant threats to real-world applications, such as face recognition [27, 37], autonomous driving [16, 44].
@@
Although many adversarial defenses against patch attacks have been proposed in the past years, the defense performance is not satisfactory, which cannot meet the demands of the safety and reliability of real-world applications. Some methods employ adversarial training [35, 25] and certified defenses [10, 4], which are only effective against specific attacks but generalize poorly to other forms of patch attacks in the real world [24]. Another category of patch defense is based on pre-processing techniques [11, 23, 40, 20], which usually destroy the patterns of adversarial patches by image completion or smoothing. However, these methods can hardly restore the images with high fidelity, leading to visual artifacts of the reconstructed images that impact recognition. They can also be evaded by stronger adaptive attacks due to gradient obfuscation [1].
@@
Figure 1: The intriguing property of the diffusion model. We perform a diffusion model multiple times on the given adversarial image, and find the differences between any two various denoised images are pronounced within the adversarial patch regions, which can be used to estimate the mask for this adversarial patch.
Recently, diffusion models [28, 14] have emerged as a powerful family of generative models, and have been successfully applied to improving adversarial robustness by purifying the input data [24, 32, 36]. By diffusing the adversarial examples with Gaussian noises and recovering the original inputs through the reverse denoising process of diffusion models, the downstream classifiers can correctly recognize the denoised images with high robustness. Our initial intuition is to explore whether diffusion purification can defend against patch attacks. However, we observe that the patch region in the adversarial image can hardly be denoised towards the clean image, while the output also differs from the raw patch. This suggests the potential to adopt diffusion models to detect the patch region, as shown in Fig. 1.
@@
Based on the observation, we propose DIFFender, a novel defense method against adversarial patch attacks with pre-trained diffusion models. DIFFender localizes the region of the adversarial patch by comparing the differences between various denoised images and then recovers the identified patch region in the image while preserving the integrity of the underlying content. Importantly, these two stages are carefully guided by a unified diffusion model, thus we can utilize the close interaction between them to improve the whole defense performance (i.e., an accurate localization will promote the following restoration, and a perfect restoration will help evaluate the performance of localization step in return). Specifically, we incorporate a text-guided diffusion model such that DIFFender can localize and recover the adversarial patches more accurately with textual prompts. Moreover, we design a few-shot prompt-tuning algorithm to facilitate simple and efficient tuning, enabling the pre-trained diffusion model to easily adapt to the adversarial defense task for improved robustness. The pipeline of DIFFender is illustrated in Fig. 2.
In summary, our contributions are as follows:
  • We discover the intriguing property of the diffusion model, that is, disparities exist in the alteration magnitude between the adversarial patch and background regions during the diffusion denoising process, where the difference between any two various denoised images can be further used to identify the adversarial patch.
  • Arising from the keen observations, we utilize a unified diffusion model throughout the entire process to localize and restore adversarial patches. Furthermore, we design an effective prompt-tuning algorithm that empowers the diffusion model to seamlessly adapt to the defense task. To the best of our knowledge, this is the first work to defend against patch attacks based on the diffusion model.
@@
  • We conduct extensive experiments on image classification, face recognition, and further in the physical world, demonstrating that DIFFender exhibits superior robustness under adversarial patch attacks. The results indicate that DIFFender can also generalize well to various scenarios, diverse classifiers, and multiple attack methods.
@@ 2 Related work
Adversarial attacks. Deep neural networks (DNNs) can be misled to produce erroneous outputs by introducing small perturbations to input examples. Most adversarial attacks [9, 22, 21, 6] typically induce misclassification or detection errors by adding small perturbations to the pixels of input examples. However, while these methods can effectively generate adversarial examples in the digital world, they lack practicality in the real world. On the other hand, adversarial patch attacks aim to deceive models by applying a pattern or a sticker to a localized region of the object, which are more realizable in the physical world [2, 17, 18, 33, 42].
Adversarial defenses. With the development of attacks, various defense methods have been proposed. However, most existing defenses primarily focus on global perturbations with $\ell_{p}$ norm constraints, including former diffusion-based defenses [24, 32, 36], and defenses against patch attacks have not been extensively studied. Despite the effectiveness of adversarial training [35, 25] and certified defenses [10, 4] against specific attacks, they have limited generalization to other patch attacks.
Therefore, most studies focus on pre-processing defenses. Digital Watermarking [11] utilizes saliency maps to detect adversarial regions and employs erosion operations to remove small holes. Local Gradient Smoothing [23] performs gradient smoothing on regions with high gradient amplitudes, taking into account the high-frequency noise introduced by patch attacks. Feature Normalization and Clipping [40] involves gradient clipping operations on images to reduce informative class evidence based on knowledge of the network structure. Jedi [31] utilizes entropy to obtain masks. However, these methods can hardly reconstruct the original image and can be evaded by adaptive attacks [1]. In contrast, we propose to leverage pre-trained diffusion models to better localize and restore the adversarial patches.
Figure 2: Pipeline of DIFFender. DIFFender leverages a unified diffusion model to jointly guide the localization and restoration of adversarial attacks, and combines a prompt-tuning module to facilitate efficient tuning.
@@ 3 Methodology
In this section, we first give the intriguing property of diffusion models to localizing patch regions, and then introduce the whole framework of our DIFFender.
3.1 Observation
@@
Figure 3: Restoration results denoised by diffusion model at noise ratios $t^{*}=0.15/0.5/0.7/0.9$. The patch cannot be removed with small ratios (see $t^{*}=0.15/0.5$), but the global structure gets lost with large ratios (see $t^{*}=0.7/0.9$).
Figure 4: We test the difference results on 100 images from ImageNet and find that differences between any two various denoised images become more pronounced within the adversarial patch regions.
@@
DiffPure [24] is a recent method that utilizes diffusion models to remove the imperceptible perturbations on the images by input purification. It first adds Gaussian noise with a noise ratio $t^{*}$ to the adversarial image and then denoises the image with the backward process of diffusion models. We first examine whether DiffPure can be applied to defend against patch attacks. As illustrated in Fig. 3, DiffPure is ineffective against patch attacks. We utilize diffusion model to purify adversarial patches at different noise ratios $t^{*}$ (parameterized within the range of 0 to 1), and find that it fails to eliminate the threat posed by adversarial patches while preserving the image semantics. This is due to the trade-off between purifying the adversarial perturbations (with a larger $t^{*}$) and preserving the image semantics (with a smaller $t^{*}$), making it impossible to find an appropriate noise ratio that can defend against adversarial patches.
@@
Although the results indicate that DiffPure is not applicable to adversarial patches, we also find that when the noise ratio is about $0.5$, the patch region in the adversarial image can hardly be denoised, but other regions outside the adversarial patch can be well-preserved. It indicates that we can calculate the difference between various denoised images to identify the region of adversarial patches. Several examples of this phenomenon are given in Fig. 4. The reason behind this phenomenon may be that adversarial patches are often meticulously crafted perturbations with a complexity far exceeding the noise present in real image datasets. Alternatively, it could be some meaningful sticker placed in an inappropriate location, signifying that the patch is out of context within the scenario. Diffusion models are trained to learn the probability distribution of real images, thus they struggle to fully adapt to the distribution of adversarial examples, leading to inaccurate estimations.
Due to the crucial observation, which can resolve the trade-off between purifying adversarial attacks and preserving image semantics when utilizing the diffusion model, we are now able to employ a single diffusion model, which first locates the patch attack and subsequently restores it.
@@ 3.2 DIFFender
Patch localization. DIFFender first performs accurate patch localization based on the above observation of the diffusion model. Given the adversarial image $\mathbf{x}_{adv}$, we first add Gaussian noise to create a noisy image $\mathbf{x}_{t}$ with noise ratio 0.5. Next, we apply a text-guided diffusion model to obtain a denoised image $\mathbf{x}_{p}$ from $\mathbf{x}_{t}$ with a textual prompt $prompt_{L}$, and $\mathbf{x}_{e}$ with empty text. We can estimate the mask region $\hat{\mathbf{M}}$ by taking the difference between the denoised images $\mathbf{x}_{p}$ and $\mathbf{x}_{e}$. However, the diffusion model incurs a significant time cost due to the time steps $T$ required. To address this issue, we directly predict the image $\mathbf{x}_{0}$ from the noisy image $\mathbf{x}_{t}$ with only one step, which saves $T$ times the processing time, based on the specific diffusion model.
Although the one-step predicted results often exhibit discrepancies and increased blurriness compared to the original image, the differences between one-step predictions still align with our observations. In practice, we perform one-step denoising twice, obtaining two results: $\mathbf{x}_{a}$, the one guided by $prompt_{L}$, and $\mathbf{x}_{b}$, the one guided by empty text to calculate the difference and binarize it, as:
@@
where we calculate the difference for $m$ times (set to 3 in the experiments) to enhance the stability, eliminate randomness, and reduce the computational complexity. The $prompt_{L}$ can be hand-designed (e.g., “adversarial”) or automatically tuned as shown in Sec. 3.3.
@@
Figure 5: The estimated difference are binarized and applied Gaussian smoothing and dilation operations to gain the refined mask.
Patch restoration. After locating the patch region, DIFFender then restores it to eliminate the adversarial effects, while also considering preserving the overall coherence and quality of the image. In particular, we combine the estimated mask $\hat{\mathbf{M}}$ and $\mathbf{x}_{adv}$ as inputs to the text-guided diffusion model with prompt $prompt_{R}$ to obtain a restored image $\mathbf{x}_{r}$. We follow the inpainting pipeline in Stable Diffusion [26] to process the mask, where a UNet is used with an additional five input channels to incorporate the estimated mask $\hat{\mathbf{M}}$. Similarly, $prompt_{R}$ can be manually set (e.g., “clean”) or automatically tuned.
To ensure the adversarial effect is removed in the images, the restoration performs the entire diffusion process, which takes several times the processing time of the localization. However, the restoration is only invoked when attacks are detected, ensuring high efficiency (See Appendix B.1).
Unified defense model. The aforementioned two stages have been meticulously integrated into one unified diffusion model (e.g., stable diffusion), driven by our critical observations. This intentional fusion allows us to harness the tight interplay between these stages, thereby enhancing the defense mechanism. As a direct consequence of our insights, we also introduced the prompt-tuning module, which encompasses the joint optimization of the entire pipeline.

### v4

@@ Abstract
Adversarial attacks, particularly patch attacks, pose significant threats to the robustness and reliability of deep learning models. Developing reliable defenses against patch attacks is crucial for real-world applications. This paper introduces DIFFender, a novel defense framework that harnesses the capabilities of a text-guided diffusion model to combat patch attacks. Central to our approach is the discovery of the Adversarial Anomaly Perception (AAP) phenomenon, which empowers the diffusion model to detect and localize adversarial patches through the analysis of distributional discrepancies. DIFFender integrates dual tasks of patch localization and restoration within a single diffusion model framework, utilizing their close interaction to enhance defense efficacy. Moreover, DIFFender utilizes vision-language pre-training coupled with an efficient few-shot prompt-tuning algorithm, which streamlines the adaptation of the pre-trained diffusion model to defense tasks, thus eliminating the need for extensive retraining. Our comprehensive evaluation spans image classification and face recognition tasks, extending to real-world scenarios, where DIFFender shows good robustness against adversarial attacks. The versatility and generalizability of DIFFender are evident across a variety of settings, classifiers, and attack methodologies, marking an advancement in adversarial patch defense strategies. Our code is available at https://github.com/kkkcx/DIFFender.
Keywords:
Adversarial Robustness Patches Attacks Diffusion Model
@@ 1 Introduction
Deep neural networks are vulnerable to adversarial examples [36, 12], in which imperceptible perturbations are intentionally added to natural examples, leading to incorrect predictions with high confidence of the model [25, 45]. Most adversarial attacks and defenses are devoted to studying the $\ell_{p}$-norm threat models [12, 3, 8, 27], which assume that the adversarial perturbations are restricted by the $\ell_{p}$ norm to be imperceptible. However, the classic $\ell_{p}$ perturbations require modification of every pixel of the images, which is typically not practical in the physical world. On the other hand, adversarial patch attacks [2, 20, 22, 39], which usually apply perturbations to a localized region of the objects, are more physically realizable. Adversarial patch attacks pose significant threats to real-world applications, such as face recognition [33, 44], autonomous driving [19, 53, 7, 21].
@@
Although many adversarial defenses against patch attacks have been proposed in the past years, the defense performance is not satisfactory, which cannot meet the demands of the safety and reliability of real-world applications. Some methods employ adversarial training [42, 31, 41] and certified defenses [13, 4], which are only effective against specific attacks but generalize poorly to other forms of patch attacks in the real world [30]. Another category of patch defense is based on pre-processing techniques [14, 29, 48, 24], which usually destroy the patterns of adversarial patches by image completion or smoothing. However, these methods can hardly restore the images with high fidelity, leading to visual artifacts of the reconstructed images that impact recognition. They can also be evaded by stronger adaptive attacks due to gradient obfuscation [1].
Recently, diffusion models [34, 17] have emerged as a powerful family of generative models, and have been successfully applied to improving adversarial robustness by purifying the input data [30, 38, 43]. Our initial intuition is to explore whether diffusion purification can defend against patch attacks. However, we find it fails to counter such attacks since it cannot remove the adversarial patches completely. In contrast, we discover the Adversarial Anomaly Perception (AAP) phenomenon, as shown in Fig. 1. The phenomenon suggests that we can calculate the difference between various denoised images to identify the region of adversarial patches. Subsequently, it facilitates targeted restoration of specific patch-affected areas. The reason behind the phenomenon may be that adversarial patches are often complexly crafted perturbations or contextually misplaced elements, significantly differing from the natural image distributions it is trained on. This phenomenon indicates progress in understanding how diffusion models can differently respond to adversarial patches and resolves the inherent trade-off between purifying the adversarial patches and preserving the image semantics.
@@
Figure 1: The intriguing phenomenon of the diffusion model. A diffusion model is performed multiple times on the given adversarial image, and the differences between any two denoised images are pronounced within the adversarial patch regions, which can be leveraged to further pinpoint the location of adversarial patches.
@@
Based on the AAP phenomenon, we propose DIFFender, a novel defense framework against adversarial patch attacks with pre-trained diffusion models. DIFFender localizes the region of the adversarial patch by comparing the differences between various denoised images and then recovers the identified patch region in the image while preserving the integrity of the underlying content. Importantly, these two stages are carefully guided by a unified diffusion model, thus we can utilize the close interaction between them to improve the whole defense (i.e., an accurate localization will promote the following restoration, and a perfect restoration will help evaluate the performance of localization step in return). Specifically, we incorporate a text-guided diffusion model such that DIFFender can localize and recover the adversarial patches more accurately with textual prompts. Moreover, we design a few-shot prompt-tuning algorithm to facilitate simple and efficient tuning, enabling the pre-trained diffusion model to easily adapt to the adversarial defense task for improved robustness. The pipeline of DIFFender is illustrated in Fig. 2. In summary, our contributions are as follows:
  • We uncover the intriguing Adversarial Anomaly Perception (AAP) phenomenon within the diffusion model, enabling it to leverage the distributional discrepancies between adversarial patches and natural images for accurate localization, thus overcoming the trade-off between purifying patches and preserving image semantics. This approach broadens the applicability of diffusion techniques, making it feasible to employ the diffusion model in countering adversarial patch attacks.
  • Arising from the AAP phenomenon, we introduce DIFFender, an innovative diffusion-based defense framework. DIFFender employs one diffusion model throughout the entire process to both localize and restore adversarial patches, leveraging vision-language pre-training to implement efficient defense. To our knowledge, DIFFender stands as the first framework to leverage the diffusion model for comprehensive defense against patch attacks, marking a notable advancement in the field.
  • Additionally, we develop an efficient prompt-tuning module and design three effective losses. The losses fine-tune the pre-trained diffusion model through the tuning process, enabling the model to co-optimize both localization and restoration modules, thereby achieving improved defense performance. This approach not only enhances the model’s adaptability but also reduces the computational overhead associated with traditional retraining methods.
@@
  • We conduct extensive experiments on image classification, face recognition, and further in the physical world, demonstrating that DIFFender effectively reduces the attack success rate even under strong adaptive attacks. The results indicate that DIFFender can also generalize well to various scenarios, diverse classifiers, and multiple attack methods.
@@ 2 Related work
Adversarial attacks. Deep neural networks (DNNs) can be misled to produce erroneous outputs [36, 12, 9] by introducing perturbations to input examples. Most adversarial attacks [12, 28, 27, 8, 26] typically induce misclassification by adding small perturbations to the pixels of input examples. However, while these methods can effectively generate adversarial examples in the digital world, they lack practicality in the real world. On the other hand, adversarial patch attacks [2, 20, 22, 39, 51] aim to deceive models by applying a pattern or a sticker to a localized region of the object, which are more realizable in the physical world.
Adversarial defenses. As attacks evolve, various defense methods have emerged. However, most existing defenses primarily focus on global perturbations with $\ell_{p}$ norm constraints, including former diffusion-based defenses [30, 38, 43], and defenses against patch attacks have not been extensively studied. Despite the effectiveness of adversarial training [42, 31, 50] and certified defenses [13, 4] against specific attacks, they have limited generalization to other patch attacks.
Therefore, most studies focus on pre-processing defenses. Digital Watermarking [14] utilizes saliency maps to detect adversarial regions and employs erosion operations to remove small holes. Local Gradient Smoothing [29] performs gradient smoothing on regions with high gradient amplitudes, taking into account the high-frequency noise introduced by patch attacks. Feature Normalization and Clipping [48] involves gradient clipping operations on images to reduce informative class evidence based on knowledge of the network structure. SAC [24] defends against patch attacks by detecting and removing patches. Jedi [37] utilizes entropy to obtain masks for patches. However, these methods can hardly reconstruct the original image and can be evaded by adaptive attacks [1]. In contrast, we introduce the AAP phenomenon and propose to leverage pre-trained diffusion models to localize and restore the adversarial patches. This method not only addresses the limitations of existing defenses but also opens new pathways for research in using diffusion models for patch defense.
Figure 2: Pipeline of DIFFender. DIFFender leverages a unified diffusion model to jointly guide the localization and restoration of adversarial patch attacks, and combines a prompt-tuning module to facilitate efficient tuning.
@@ 3 Methodology
In this section, we first introduce the discovery of the AAP phenomenon within diffusion models. Following this understanding, we outline the whole framework of DIFFender and detail the improved techniques by prompt tuning.
@@
Figure 3: Denoised results by diffusion model at different noise ratios. With small ratios ( $t^{*}=0.15/0.5$ ), the patch cannot be purified; conversely, the global structure becomes lost with large ratios ( $t^{*}=0.7/0.9$ ).
Figure 4: In the analysis of ImageNet images, we find a pronounced difference specifically within regions affected by adversarial patches. This observation provides empirical evidence supporting the AAP phenomenon.
3.1 Discovery of the AAP Phenomenon
@@
DiffPure [30] is a recent method that utilizes diffusion models to remove the imperceptible perturbations on the images by introducing Gaussian noise at a predetermined ratio $t^{*}$ (ranging from 0 to 1) to adversarial images, followed by a denoising process via the reverse dynamics of diffusion models. Our study initially aimed to assess the applicability of DiffPure against patch attacks. The empirical investigations, illustrated in Fig. 3 demonstrate the inadequacy of DiffPure in countering patch attacks. This inefficacy stems from an inherent trade-off between purifying the adversarial perturbations (with a larger $t^{*}$) and preserving the image semantics (with a smaller $t^{*}$), making it impossible to find an appropriate noise ratio that can defend against adversarial patches.
@@
In contrast, we find at a critical noise ratio of $t^{*}$, a distinct pattern emerged: while the adversarial patches exhibited resistance to denoising, they also struggled to be restored, resulting in variable outcomes. Meanwhile, the remainder of the image retained its semantic integrity unscathed. This suggests that we can calculate the difference between various denoised images to identify the region of adversarial patches. This observation, depicted through various examples in Fig. 4, leads to the Adversarial Anomaly Perception (AAP) phenomenon. The reason behind this phenomenon may be that adversarial patches are often intentionally crafted perturbations with a complexity far exceeding the noise present in real image datasets. Alternatively, it could be some meaningful sticker placed in an inappropriate location, signifying that the patch is out of context within the scenario. Diffusion models are trained to learn the probability distribution of real images, thus they struggle to fully adapt to the distribution of adversarial examples in its latent space, leading to the difference.
The discovery of AAP provides insight into understanding how diffusion models can differentially respond to adversarial patches, which empowers the diffusion model to detect and localize adversarial patches through the analysis of distributional discrepancies, and further facilitates targeted restoration of specific patch-affected areas. It resolves the inherent trade-off between purging adversarial patches and preserving image authenticity. Leveraging AAP, we propose a unified diffusion-based defense framework DIFFender, employing a single diffusion model that both locates and restores patch attacks.
@@ 3.2 DIFFender
Patch localization. DIFFender first performs accurate patch localization based on the above phenomenon of the diffusion model. Given the adversarial image $\mathbf{x}_{adv}$, we first add Gaussian noise to create a noisy image $\mathbf{x}_{t}$ with a certain noise ratio $t^{*}$ (chosen as 0.5 in the experiments). Next, inspired by [5], we apply a text-guided diffusion model to obtain a denoised image $\mathbf{x}_{p}$ from $\mathbf{x}_{t}$ with a textual prompt $prompt_{L}$, and $\mathbf{x}_{e}$ with empty text. We can estimate the mask region $\hat{\mathbf{M}}$ by taking the difference between the denoised images $\mathbf{x}_{p}$ and $\mathbf{x}_{e}$. However, the diffusion model incurs a significant time cost due to the time steps $T$ required. To address this issue, we directly predict the image $\mathbf{x}_{0}$ from the noisy image $\mathbf{x}_{t}$ with only one step, which saves $T$ times the processing time.
Although the one-step predicted results often exhibit discrepancies and increased blurriness compared to the original one, the differences between one-step predictions still align with the AAP phenomenon. In practice, we perform one-step denoising twice, obtaining two results: $\mathbf{x}_{a}$, the one guided by $prompt_{L}$, and $\mathbf{x}_{b}$, the one guided by empty text to calculate the difference and binarize it, as:
@@
where we calculate the difference for $m$ times (set to 3 in the experiments) to enhance stability and eliminate randomness. The $prompt_{L}$ can be hand-designed (e.g., "adversarial”) or automatically tuned as shown in Sec. 3.3.
@@
Figure 5: To gain the final refined mask, the estimated differences are binarized, applied Gaussian smoothing and dilation operations.
Patch restoration. After locating the patch region, DIFFender then restores it to eliminate the adversarial effects, while also considering preserving the overall coherence and quality of the image. In particular, we combine the estimated mask $\hat{\mathbf{M}}$ and $\mathbf{x}_{adv}$ as inputs to the text-guided diffusion model with prompt $prompt_{R}$ to obtain a restored image $\mathbf{x}_{r}$. We follow the inpainting pipeline in Stable Diffusion [32] to process the mask, where a UNet is used with an additional five input channels to incorporate the estimated mask $\hat{\mathbf{M}}$. Similarly, $prompt_{R}$ can be manually set (e.g., "clean”) or automatically tuned.
Unified defense model. The aforementioned two stages have been meticulously integrated into one unified diffusion model (e.g., Stable Diffusion), driven by the critical AAP phenomenon. This intentional fusion allows us to harness the tight interplay between these stages, thereby enhancing the defense mechanism. As a direct consequence of our insights, we also introduced the prompt-tuning module, which encompasses the joint optimization of the entire pipeline.